### v1

@@ IV-C Ablation Studies
IV-D Computational Cost Comparison
Figure 9: Visual examples of different methods on the DroneVehicle dataset. From left to right are Halfway Fusion [55], CIAN [12] and AR-CNN [20] and our $\textup{C}^{2}$ormer-$\textup{S}^{2}$et, respectively. The confidence threshold is set to 0.3 when visualizing these results.
IV-E Visualization of Feature Maps
To illustrate the effectiveness of ICA module, we visualize the intermediate results in Figure 7. From the first row, we find that the similarity map corresponding to the red and green points can clearly reflect the unaligned position in the other modality, which shows this similarity map indeed has the ability to localize the calibration position. From the $y_{ir}$ in the second row, we can see that the ICA module can dynamically complement the RGB features according to the intensity of the IR features, for example, the obscure features of $x_{rgb}$ are complemented in the red box. All the results demonstrate that the ICA module achieves both feature calibration and complementary fusion.
Since the MBNet [16] method is a popular method by fusing RGB-IR features in the backbone, we also visualize the feature maps (shown in Figure 8) of MBNet method and $\textup{C}^{2}$ormer respectively. From the results, we note that vehicles show more contrast with the background in the feature maps generated by $\textup{C}^{2}$ormer for all capturing times. During the daytime (first row), objects with miscalibration problem (inside the red box) have better feature representation by using $\textup{C}^{2}$ormer. Furthermore, when the objects are in the low-light condition (second and third rows), $\textup{C}^{2}$ormer can complement corresponding features with global information to enhance the feature representation compared to the MBNet method.

### v3

@@ IV-C Ablation Studies
IV-D Computational Cost Comparison
Figure 9: Detection results of our $\textup{C}^{2}$ormer in different scenarios on the DroneVehicle dataset. From top to bottom, objects are in different complex situations of occlusion, low resolution and angle changes.
Figure 10: Visual examples of different methods on the DroneVehicle dataset. From top to bottom are GroundTruth (IR and RGB), Halfway Fusion [67], CIAN [15], AR-CNN [25] and our $\textup{C}^{2}$ormer-$\textup{S}^{2}$et, respectively. The confidence threshold is set to 0.3 when visualizing these results.
IV-E Visualization Results
Intermediate results of ICA module. To illustrate the effectiveness of ICA module, we visualize the intermediate results in Figure 7. From the first row, we find that the similarity map corresponding to the red and green points can clearly reflect the unaligned position in the other modality, which shows this similarity map indeed has the ability to localize the calibration position. From the $y_{ir}$ and $y_{rgb}$ in the second row, we can observe that the ICA module can dynamically complement the RGB and IR features according to the intensity of the $x_{rgb}$ and $x_{ir}$. For example, the obscure features $x_{rgb}$ and $x_{ir}$ in the red box are complemented. The results show that the ICA module achieves both feature calibration and complementary fusion.
Visualization of feature map. Since the MBNet [19] method is a popular method by fusing RGB-IR features in the backbone, we also visualize the feature maps (shown in Figure 8) of MBNet method and $\textup{C}^{2}$ormer respectively. We note that vehicles show more contrast with the background in the feature maps generated by $\textup{C}^{2}$ormer for all capturing times. During the daytime (first row), objects with miscalibration problem (inside the red box) have better feature representation by using $\textup{C}^{2}$ormer. Furthermore, when the objects are in the low-light condition (second and third rows), $\textup{C}^{2}$ormer can complement corresponding features with global information to enhance the feature representation compared to the MBNet method.
Detection results in complex scenarios. To illustrate the detection capabilities of our model in complex scenarios, we visualize the results of the model in different scenarios. As shown in Figure 9, the objects in these images are in three complex situations: occlusion, low resolution and angle change. Through these visual results, we find that $\textup{C}^{2}$ormer is robust to the above situations, which verifies its effectiveness.